OpenStudy (anonymous):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –3 7 –2 5 Line 2 x y 0 5 2 1 Line 1: y = x + Line 2: y = x + @dtan5457

OpenStudy (dtan5457):

easy peezy give me 2 min

OpenStudy (dtan5457):

Ok first line i'll show you. slope intercept form is y=mx+b it's important to understand m is the slope, and b is the y intercept you have the points (-3,7) (-2,5) to put in slope intercept..you first must need the slope y2-y1/x2-x1 plug those in and you find that the slope is -2 perfect, but now you need to plug in any point that they gave you into the slope intercept so like im gonna use -3,7 into y=mx+b plug everything in to get 7=-2(-3)+b solve for b and it's easily 7=6+b b=1 now just finalize your equation to y=-2x+1<<<answer for 1st line

OpenStudy (dtan5457):

wanna try the 2nd one yourself?

OpenStudy (dtan5457):

it's fun once you learn it :D

OpenStudy (anonymous):

give me one minute im working it out

OpenStudy (dtan5457):

alright, take your time. i'm doing my own homework too so i won't be going anywhere else

OpenStudy (dtan5457):

Did you get it?

OpenStudy (anonymous):

y2-y1/x2-x1= 1-5/2-0 =-4/2 slope is -4/2 y=mx+b 5=-4/2*0+b 5=0+b b=5 I got up to here now im lost??

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Yup, you got everything right here so far. Put the -4/2 slope as -2, (easier to read) now, you know that the slope is -2. the y intercept is 5. simply plug it in to the y=mx+b formula y=-2x+5 m=slope b=y intercept easy as that

OpenStudy (dtan5457):

questions?

OpenStudy (anonymous):

one more What is the solution to the system? Line 1 x y –2 –4 –1 –1 Line 2 x y 2 3 3 1

OpenStudy (dtan5457):

ok. another few min.

OpenStudy (dtan5457):

this will be long. lol.

OpenStudy (dtan5457):

you need to make 2 linear equations from each set of points. first line: (-2,-4) (-1,-1) from what we learned before, to make the equation we need the slope. i won't bother showing work as i think you got that by now. slope=3x ill use point slope formula to get the equation y-y1=m(x-x1) i use the point -2,-4 you get y+4=3(x+2) simplify that to y=3x+2<ur first equation for the first set 2nd set, same stuff (2,3) (3,1) slope, point slope or intercept, all that stuff slope:-2x y-3=-2(x-2) y=-2x+7<<<<ur 2nd equation heres the tricky part, equate these equations together -2x+7=3x+2 solve for x, and you get -5x=-5, or x=1 plug x=1 to any of those lines to get y y=3(1)+2, y=5 your solution=(1,5)